OpenStudy (anonymous):

7.2-12). The following problem involves some factoring by grouping. Remember, before you can divide out factors common to the numerators and denominators of a product, you must factor completely. 2x^2+4x x^2+3x+xy+3y _________ . _______________ x^2-y^2 x^2+5+6

OpenStudy (watchmath):

\(\frac{2x(x+2)}{(x+y)(x-y)}\cdot\frac{(x+y)(x+3)}{(x+3)(x+2)}=\frac{2x}{x-y}\)